OpenStudy (anonymous):

what is (sqrt(3)/2) / (1- 1/2)

OpenStudy (anonymous):

no wolframs please:)

OpenStudy (diyadiya):

\[\large \frac{\frac{\sqrt{3}}{2}}{1- \frac{1}{2}}=\large \frac{\frac{\sqrt{3}}{2}}{ \frac{1}{2}}= \sqrt{3}\]

OpenStudy (anonymous):